# Manifolds with kinks and the asymptotic behavior of the graph Laplacian operator with Gaussian kernel

### v3

@@ 1. Introduction
The connection between the graph Laplacian and the Laplace–Beltrami operator is a central theme in geometric data analysis and has attracted substantial interest from both applied and theoretical communities. On the one hand, from a theoretical perspective, this relationship offers a discretization framework for studying differential operators on manifolds via point clouds or sampled data [LiShiSun, HarlimSanzYang, YanJiangHarlim, JiaoYanHarlimLu]. On the other hand, from an applied viewpoint, this connection underpins a wide array of algorithms in machine learning and signal processing, including manifold learning techniques such as Laplacian eigenmaps [BelkinNiyogi2003, BelkinNiyogi2005] or Diffusion Maps [CoifmanLafon2006, SingerWu]. Understanding the convergence of the graph Laplacian to the Laplace–Beltrami operator — notably the role of scaling, sampling density, and boundary behavior — remains crucial in ensuring that discrete approximations faithfully capture the geometry and analysis of the underlying continuous space. This paper aims at tackling this question on a class of smooth Riemannian manifolds having possibly singular boundary.
@@
After the work of various authors, e.g. [Hein2005, BelkinNiyogi2005], it is by now known that when $M$ is a smooth $d$-dimensional submanifold of an Euclidean space $\mathbb{R}^{D}$, for any sufficiently regular $f$ and $p$, the extrinsic Gaussian operator satisfies
(1) $L_{t}f(x)\to c_{d}\,\Delta_{M,p}f(x)\qquad\text{ as $t\to 0$}$
at any interior point $x\in M$, where $c_{d}$ is a dimensional constant and $\Delta_{M,p}$ is the weighted Laplace operator given by
$\Delta_{M,p}f(x):=p(x)\Delta_{M}{f}(x)+\frac{1}{2}\nabla{p}(x)\cdot\nabla{f}(x),$
with $\Delta_{M}$ the classical Laplace–Beltrami operator of $M$. In [belkin+] (see also [AnderssonAvelin]), this pointwise result for the extrinsic Gaussian operator was extended to three types of singular subsets of $\mathbb{R}^{D}$, namely submanifolds with boundaries, intersections of them, and submanifolds with edge-type singularities. For instance, if $x$ belongs to the non-empty boundary of some smooth submanifold $M\subset\mathbb{R}^{D}$, then
(2) $L_{t}f(x)=\frac{c^{\prime}_{d}}{\sqrt{t}}p(x)\partial_{\nu}f(x)+o\left(\frac{1}{\sqrt{t}}\right)\qquad\text{ as $t\downarrow 0$,}$
where $\partial_{\nu}f(x)$ is the inner normal derivative of $f$ at $x$ and $c_{d}^{\prime}$ is another dimensional constant.
Our main results extend both (1) and (2) in the setting of smooth Riemannian manifolds with kinks. These spaces, which we introduce in Section 3, might be understood as manifolds with possibly non-empty and irregular boundary. To distinguish these possibly non-smooth boundaries with classical smooth ones, we use the terminology border : a smooth manifold with kinks $M$ thus divides into an interior part $intM$ and a border part $\partial M$. Manifolds with kinks notably include manifolds without boundary, manifolds with smooth boundary, and manifolds with corners as considered e.g. in [Cerf, Douady, DouadyHerault, Michor, Melrose, Joyce]. Other relevant spaces like cones, pyramids, even cusps, fall into the framework of manifolds with kinks while they are not manifolds with corners.
@@
For any positive integer $\ell$ we set $c_{\ell}:=\Gamma((\ell+1)/2)/2$ where $\Gamma$ is the classical Gamma function. Then our first main result concerning the intrinsic Gaussian operator writes as follows.
@@
as $t\downarrow 0$, for any $f\in\mathcal{C}^{3}(M)\cap L^{1}(M,p\,\mathrm{vol}_{g})$, where $\mathrm{Err}(t)=o(1)$ as $t\downarrow 0$, and
@@
$A_{g}f(x)$ $:=\int_{S^{g}I_{x}M}\mathop{}\!\mathrm{d}_{x}^{(2)}f\left(\theta^{(2)}\right)\mathop{}\!\mathrm{d}\sigma(\theta),$
$[p,f]_{g}(x)$ $:=\frac{1}{2}\int_{S^{g}I_{x}M}\mathop{}\!\mathrm{d}_{x}f(\theta)\mathop{}\!\mathrm{d}_{x}p(\theta)\mathop{}\!\mathrm{d}\sigma(\theta).$
@@
    If $x$ is an interior point, then $\mathrm{Err}(t)=0$ for any small enough $t>0$.
@@
    (3) $\mathrm{Err}(t)\leq\,4\Gamma(d+2)\,p(x)\,(\|d_{x}f\|_{g_{x}}+\|d_{x}^{(2)}f\|_{g_{x}})\,SC(x).$
    where $SC(x)$ is the total slicewise curvature of $\partial\Omega$ at $x$, defined in (38). In particular, if the total slicewise curvature vector is zero, then $\mathrm{Err}(t)=0$ for any small enough $t>0$.
@@
$A_{M}\tilde{f}(0_{d}):=\int_{S^{B}_{0_{d}}\Omega}\mathop{}\!\mathrm{d}_{x}^{(2)}\tilde{f}\left(\theta^{(2)}\right)\mathop{}\!\mathrm{d}\sigma(\theta),\qquad[\tilde{p},\tilde{f}]_{M}(0_{d}):=\frac{1}{2}\int_{S^{B}_{0_{d}}\Omega}\mathop{}\!\mathrm{d}_{x}\tilde{f}(\theta)\mathop{}\!\mathrm{d}_{x}\tilde{p}(\theta)\mathop{}\!\mathrm{d}\sigma(\theta).$
@@
It is worth pointing out that in the previous context, the inward tangent sector $I_{x}M$ coincides with the Bouligand tangent cone of $\Omega$ at $0_{d}$, so that Theorem 1.2 matches up with Theorem 1.1. Let us also stress out that, while bounded geometry of the boundary was assumed in [HAL2007], we do not make any assumption of that kind here.
The recent preprint [ART] provides a result that goes in the direction of Theorem 1.2, but with less details on the differential operators involved there. Our approach allows for a neat identification of the latter in several cases, including the following ones.
  • If $x$ is an interior point, then $I_{x}M$ coincides with $T_{x}M$. The symmetry of $T_{x}M$ with respect to the antipodal map $\theta\mapsto-\theta$ implies that $v_{g}(x)=0$, while $p(x)A_{g}f(x)+[p,f]_{g}(x)$ reduces to the weighted Laplacian $p(x)\Delta_{g}(x)+\frac{1}{2}g_{x}(\nabla^{g}{p}(x),\nabla^{g}{f}(x))$ where $\Delta_{g}$ is the Laplace–Beltrami operator associated with the metric $g$. This is (1).
  • If $x$ is a $\mathcal{C}^{1}$ boundary point, we obtain
    $I_{x}M=\{v\in T_{x}M:g(\partial_{\nu}^{g}f(x),v)\geq 0\}$
    where $\partial_{\nu}^{g}f(x)$ is the inner normal derivative of $f$ at $x$, hence $\partial_{v_{g}(x)}f(x)$ coincides with $\partial_{\nu}^{g}f(x)$ and we recover (2).
  • If $x$ is an essential corner of depth $k$, meaning that $M$ is locally modeled on
@@
    where each $\partial_{i}f(x)$ corresponds to differentiation along a corresponding half-line direction of $\mathbb{R}^{n}_{k}$, and $\mathop{}\!\mathrm{d}_{x}f(v_{g}(x))=\sum_{i}\partial_{i}f(x)$. This seems to be new.
  • Other cases where $x$ is not even an essential corner, like tips of cones and pyramids, can be investigated in the same way.
Theorems 1.1 and 1.2 open the door to a natural Neumann boundary condition on (sub)manifolds with kinks, that would be $\partial_{v_{g}(x)}f(x)=0$ for any $x\in\partial M$. A natural problem to tackle after that is the spectral convergence of the Graph Laplacian with Gaussian kernel generated from finitely many samples $\{X_{1},\ldots,X_{n}\}$ to the suitable Neumann Laplacian as sample size goes to infinity. We refer to [BelkinNiyogi_conv, SingerWu2, TGHS] for related results on manifolds without and with boundary, see also [DT] for a preliminary investigation with the symmetrized AMV kernel.
@@
We obtain Theorem 1.3 in Section 7 through concentration estimates. Such estimates — studied in e.g. [Hein2005, belkin+, AnderssonAvelin] — quantify the error in approximating the Gaussian operator by the graph Laplace operator. Our results generalize this line of work in two key directions :
@@
    We work in the setting of Riemannian manifolds with kinks, whose singularities can be significantly more severe than those of manifolds with boundary considered in [belkin+].
@@
Acknowledgments.
The authors are funded by the Research Foundation – Flanders (FWO) via the Odysseus II programme no. G0DBZ23N. They both thank Dominic Joyce for helpful answers concerning manifolds with corners, Iosif Pinelis who suggested the proof of (7), and Laurent Bessières for a comment on the first version of the paper.
@@ 2.1. Boundary points
For an open set $\Omega\subset\mathbb{R}^{d}$ with a non-empty boundary $\partial\Omega$, we will call boundary points of $\Omega$ the elements in $\partial\Omega$. Inspired by [Evans, p.626], we define the regularity of boundary points as follows.
@@ 2.2. Diffeomorphisms and inward sectors
For $k$ a positive integer, we shall use the following notion of $\mathcal{C}^{k}$ diffeomorphism between possibly non-open subsets of $\mathbb{R}^{d}$, see e.g. [Melrose, Section 1.5] for details.
@@ 2.3. Boundary cones of open Euclidean subsets
In this section, we provide classical notions from convex analysis, see e.g. [RW] for more details. We recall that $C\subset\mathbb{R}^{d}$ is called a cone if it is invariant under multiplication by a positive number, i.e. if $\lambda v\in C$ for any $v\in C$ and $\lambda>0.$
@@
The Bouligand tangent cone is also known as tangent cone or contingent cone in the literature. It is a closed set. One can easily check that our definition is equivalent to the one given in [RW, Def. 6.1]. Note also that
@@ 3.3.1. Tangent space
From now on, we work only with smooth manifolds with kinks, i.e. those manifolds with kinks for which the transition maps are all $\mathcal{C}^{\infty}$. This is to make sure that the tangent space defined below as the space of smooth derivations (i.e. linear forms on the space of smooth functions) is finite dimensional. Indeed, for any $\mathcal{C}^{k}$ manifold with $k<+\infty$, the space of $\mathcal{C}^{k}$ derivations is infinite dimensional : see e.g. [AMR, Theorem 4.2.41].
In order to introduce smooth derivations on manifolds with kinks, we must first define what a smooth function is. This is what we do in the next definition, largely inspired by the context of manifolds with corners (see e.g. [Joyce, Definition 2.2]).
@@
If $x\in\partial M$ and $f\circ\phi^{-1}:\phi(\bar{U})\cap\mathbb{B}_{\rho}^{d}\to\mathbb{R}$ admits Taylor polynomials of arbitrary high order, then the existence of a smooth extension of $f\circ\phi^{-1}$ to $\mathbb{R}^{d}$ is ensured by Whitney’s extension theorem [Whitney, Theorem I], see also [Stein, VI.2.]. Note that in the setting of manifolds with corners, the latter upgrades into the well-posedness of Seeley’s linear extension operator [Seeley], see [Melrose, Section 1.4].
@@
We can now define the differential of a smooth function between smooth manifolds with kinks. There is no change with the case of classical manifolds, see e.g. [Lee, p.55].
@@
When $x$ is an interior point, a $\mathcal{C}^{1}$ boundary point or an essential corner, this is already known, see [Joyce]. Let us then assume that $x$ is an essential kink which is not an essential corner. Consider a local chart $(U,\phi)$ centered at $x$. Like for manifolds without boundary, we obtain from the chain rule applied to the identities $\phi\circ\phi^{-1}=\mathrm{id}_{\phi(U)}$ and $\phi^{-1}\circ\phi=\mathrm{id}_{U}$ that the differential $\mathop{}\!\mathrm{d}_{x}\phi:T_{x}M\to T_{0_{d}}\phi(\bar{U})$ is a linear isomorphism with inverse $\mathop{}\!\mathrm{d}_{0_{d}}\phi^{-1}$. But the space of derivations at the $\mathcal{C}^{0}$ boundary point $0_{d}$ of $\phi(U)$ coincides with $\mathbb{R}^{n}$, since a basis of this space is given by the classical partial differential operators $\partial\cdot/\partial x_{1},\ldots,\,\partial\cdot/\partial x_{n}$ defined on the open set $\phi(U)$ and naturally extended to any open set $V\subset\mathbb{R}^{d}$ containing $\overline{\phi(U)}$. ∎
@@ 3.3.2. Inward sector
Let us now introduce the notion of inward tangent sector for smooth manifolds with kinks. This is analogous to the inward tangent sector of Euclidean domains as discussed in Section 2.2. Our definition builds upon the classical characterization of the tangent space in terms of initial velocities of curves, which holds true for manifolds without/with boundary (see e.g. [Lee, p. 68–70]). For manifolds with boundary, one can identify inward tangent vectors at boundary points by specifying the domain of the curves we choose. This is how we came up with the following natural definition.

### v4

@@ Abstract.
Keywords. Graph Laplacian, Manifolds with non-smooth boundary, Bouligand tangent cone, nonlinear dimensional reduction, concentration estimates.
MSC Numbers. 58J32, 49J52, 68R10.
*Corresponding author.
@@ 1. Introduction
The connection between the graph Laplacian and the Laplace–Beltrami operator is a central theme in geometric data analysis and has attracted substantial interest from both applied and theoretical communities. On the one hand, from a theoretical perspective, this relationship offers a discretization framework for studying differential operators on manifolds via point clouds or sampled data [22, 16, 35, 19]. On the other hand, from an applied viewpoint, this connection underpins a wide array of algorithms in machine learning and signal processing, including manifold learning techniques such as Laplacian eigenmaps [4, 5] or Diffusion Maps [9, 30]. Understanding the convergence of the graph Laplacian to the Laplace–Beltrami operator — notably the role of scaling, sampling density, and boundary behavior — remains crucial in ensuring that discrete approximations faithfully capture the geometry and analysis of the underlying continuous space. This paper aims at tackling this question on a class of smooth Riemannian manifolds having possibly singular boundary.
@@
After the work of various authors, e.g. [18, 5], it is by now known that when $M$ is a smooth $d$-dimensional submanifold of an Euclidean space $\mathbb{R}^{D}$, for any sufficiently regular $f$ and $p$, the extrinsic Gaussian operator satisfies
(1) $L_{t}f(x)\to-\frac{\pi^{d/2}}{4}\,\Delta_{M,p}f(x)\qquad\text{ as $t\to 0$}$
at any interior point $x\in M$, where $\Delta_{M,p}$ is the weighted Laplace operator given by
$\Delta_{M,p}f(x):=p(x)\Delta_{M}{f}(x)+2\nabla{p}(x)\cdot\nabla{f}(x),$
with $\Delta_{M}$ the classical Laplace–Beltrami operator of $M$. In [7] (see also [3]), this pointwise result for the extrinsic Gaussian operator was extended to three types of singular subsets of $\mathbb{R}^{D}$, namely submanifolds with boundaries, intersections of them, and submanifolds with edge-type singularities. For instance, if $x$ belongs to the non-empty boundary of some smooth submanifold $M\subset\mathbb{R}^{D}$, then
(2) $L_{t}f(x)=-\frac{\pi^{(d-1)/2}}{2\sqrt{t}}p(x)\partial_{\nu}f(x)+o\left(\frac{1}{\sqrt{t}}\right)\qquad\text{ as $t\downarrow 0$,}$
where $\partial_{\nu}f(x)$ is the inner normal derivative of $f$ at $x$. Note that the case of isolated interior singularities has been recently addressed in [26].
Our main results extend both (1) and (2) in the setting of smooth Riemannian manifolds with kinks. These spaces, which we introduce in Section 3, might be understood as manifolds with possibly non-empty and irregular boundary. To distinguish these possibly non-smooth boundaries with classical smooth ones, we use the terminology border : a smooth manifold with kinks $M$ thus divides into an interior part $intM$ and a border part $\partial M$. Manifolds with kinks notably include manifolds without boundary, manifolds with smooth boundary, and manifolds with corners as considered e.g. in [8, 12, 11, 24, 23, 20]. Other relevant spaces like cones, pyramids, even cusps, fall into the framework of manifolds with kinks while they are not manifolds with corners.
@@
For any $\ell\in\mathbb{N}$, set
$c_{\ell}=\frac{1}{2}\Gamma((\ell+1)/2)$
where $\Gamma$ denotes the classical Gamma function. Our first main result concerning the intrinsic Gaussian operator writes as follows.
@@
as $t\downarrow 0$, for any $f\in\mathcal{C}^{3}(M)\cap L^{1}(M,p\,\mathrm{vol}_{g})$, $\mathrm{Err}(t)=o(1)$ as $t\downarrow 0$, and
@@
$A_{g}f(x)$ $:=\frac{1}{2}\int_{S^{g}I_{x}M}\mathop{}\!\mathrm{d}_{x}^{(2)}f\left(\theta^{(2)}\right)\mathop{}\!\mathrm{d}\sigma(\theta),$
$[p,f]_{g}(x)$ $:=\int_{S^{g}I_{x}M}\mathop{}\!\mathrm{d}_{x}f(\theta)\mathop{}\!\mathrm{d}_{x}p(\theta)\mathop{}\!\mathrm{d}\sigma(\theta).$
@@
    If $x$ is an interior point, then as $t\downarrow 0$,
    $L_{t}f(x)=-\frac{\pi^{d/2}}{4}\,\Delta_{M,p}f(x)+O(\sqrt{t})+O(t^{-d}e^{-t^{2\eta-1}}).$
  2. (2)
    If $x$ is a $\mathcal{C}^{1}$ boundary point, then as $t\downarrow 0$,
    (3) $L_{t}f(x)=-\frac{\pi^{(d-1)/2}}{2\sqrt{t}}\bigg(p(x)\partial_{\nu}f(x)+\mathrm{Err}(t)\bigg)-c_{d+1}\bigg(p(x)A_{g}f(x)+[p,f]_{g}(x)\bigg)+O(\sqrt{t})+O(t^{-d}e^{-t^{2\eta-1}})$
@@
    (4) $\mathrm{Err}(t)\leq\,\sqrt{t}\,C_{d}\,p(x)\,\|d_{x}f\|_{g_{x}}SC(x).$
    where $C_{d}=8\Gamma(d+2)$ and $SC(x)$ is the total slicewise curvature of $\partial\Omega$ at $x$ defined in (39). In particular, if the total slicewise curvature vector is zero, then $\mathrm{Err}(t)=0$ for any small enough $t>0$.
@@
$A_{M}\tilde{f}(0_{d}):=\frac{1}{2}\int_{S^{B}_{0_{d}}\Omega}\mathop{}\!\mathrm{d}_{x}^{(2)}\tilde{f}\left(\theta^{(2)}\right)\mathop{}\!\mathrm{d}\sigma(\theta),\qquad[\tilde{p},\tilde{f}]_{M}(0_{d}):=\int_{S^{B}_{0_{d}}\Omega}\mathop{}\!\mathrm{d}_{x}\tilde{f}(\theta)\mathop{}\!\mathrm{d}_{x}\tilde{p}(\theta)\mathop{}\!\mathrm{d}\sigma(\theta).$
@@
It is worth pointing out that in the previous context, the inward tangent sector $I_{x}M$ coincides with the Bouligand tangent cone of $\Omega$ at $0_{d}$, so that Theorem 1.2 matches up with Theorem 1.1. Let us also stress out that, while bounded geometry of the boundary was assumed in [17], we do not make any assumption of that kind here.
The recent preprint [2] provides a result that goes in the direction of Theorem 1.2, but with less details on the differential operators involved there.
Let us also point out that, if $x$ is a corner of depth $k$, meaning that $M$ is locally modeled on
@@
where each $\partial_{i}f(x)$ corresponds to differentiation along a corresponding half-line direction of $\mathbb{R}^{n}_{k}$, and $\mathop{}\!\mathrm{d}_{x}f(v_{g}(x))=\sum_{i}\partial_{i}f(x)$.
Theorems 1.1 and 1.2 open the door to a natural Neumann boundary condition on (sub)manifolds with kinks, that would be $\partial_{v_{g}(x)}f(x)=0$ for any $x\in\partial M$. A natural problem to tackle after that is the spectral convergence of the Graph Laplacian with Gaussian kernel generated from finitely many samples $\{X_{1},\ldots,X_{n}\}$ to the suitable Neumann Laplacian as sample size goes to infinity. We refer to [6, 31, 14] for related results on manifolds without and with boundary, see also [10] for a preliminary investigation with the symmetrized AMV kernel.
@@
We obtain Theorem 1.3 in Section 7 through concentration estimates. Such estimates — studied in e.g. [18, 7, 3] — quantify the error in approximating the Gaussian operator by the graph Laplace operator. Our results generalize this line of work in two key directions :
@@
    We work in the setting of Riemannian manifolds with kinks, whose singularities can be significantly more severe than those of manifolds with boundary considered in [7].
@@ 2.1. Boundary points
For an open set $\Omega\subset\mathbb{R}^{d}$ with a non-empty boundary $\partial\Omega$, we will call boundary points of $\Omega$ the elements in $\partial\Omega$. Inspired by [13, p.626], we define the regularity of boundary points as follows.
@@ 2.2. Diffeomorphisms and inward sectors
For $k$ a positive integer, we shall use the following notion of $\mathcal{C}^{k}$ diffeomorphism between possibly non-open subsets of $\mathbb{R}^{d}$, see e.g. [23, Section 1.5] for details.
@@ 2.3. Boundary cones of open Euclidean subsets
In this section, we provide classical notions from convex analysis, see e.g. [28] for more details. We recall that $C\subset\mathbb{R}^{d}$ is called a cone if it is invariant under multiplication by a positive number, i.e. if $\lambda v\in C$ for any $v\in C$ and $\lambda>0.$
@@
The Bouligand tangent cone is also known as tangent cone or contingent cone in the literature. It is a closed set. One can easily check that our definition is equivalent to the one given in [28, Def. 6.1]. Note also that
@@ 3.3.1. Tangent space
From now on, we work only with smooth manifolds with kinks, i.e. those manifolds with kinks for which the transition maps are all $\mathcal{C}^{\infty}$. This is to make sure that the tangent space defined below as the space of smooth derivations (i.e. linear forms on the space of smooth functions) is finite dimensional. Indeed, for any $\mathcal{C}^{k}$ manifold with $k<+\infty$, the space of $\mathcal{C}^{k}$ derivations is infinite dimensional : see e.g. [1, Theorem 4.2.41].
In order to introduce smooth derivations on manifolds with kinks, we must first define what a smooth function is. This is what we do in the next definition, largely inspired by the context of manifolds with corners (see e.g. [20, Definition 2.2]).
@@
If $x\in\partial M$ and $f\circ\phi^{-1}:\phi(\bar{U})\cap\mathbb{B}_{\rho}^{d}\to\mathbb{R}$ admits Taylor polynomials of arbitrary high order, then the existence of a smooth extension of $f\circ\phi^{-1}$ to $\mathbb{R}^{d}$ is ensured by Whitney’s extension theorem [34, Theorem I], see also [32, VI.2.]. Note that in the setting of manifolds with corners, the latter upgrades into the well-posedness of Seeley’s linear extension operator [29], see [23, Section 1.4].
@@
We can now define the differential of a smooth function between smooth manifolds with kinks. There is no change with the case of classical manifolds, see e.g. [21, p.55].
@@
When $x$ is an interior point, a $\mathcal{C}^{1}$ boundary point or an essential corner, this is already known, see [20]. Let us then assume that $x$ is an essential kink which is not an essential corner. Consider a local chart $(U,\phi)$ centered at $x$. Like for manifolds without boundary, we obtain from the chain rule applied to the identities $\phi\circ\phi^{-1}=\mathrm{id}_{\phi(U)}$ and $\phi^{-1}\circ\phi=\mathrm{id}_{U}$ that the differential $\mathop{}\!\mathrm{d}_{x}\phi:T_{x}M\to T_{0_{d}}\phi(\bar{U})$ is a linear isomorphism with inverse $\mathop{}\!\mathrm{d}_{0_{d}}\phi^{-1}$. But the space of derivations at the $\mathcal{C}^{0}$ boundary point $0_{d}$ of $\phi(U)$ coincides with $\mathbb{R}^{n}$, since a basis of this space is given by the classical partial differential operators $\partial\cdot/\partial x_{1},\ldots,\,\partial\cdot/\partial x_{n}$ defined on the open set $\phi(U)$ and naturally extended to any open set $V\subset\mathbb{R}^{d}$ containing $\overline{\phi(U)}$. ∎
@@ 3.3.2. Inward sector
Let us now introduce the notion of inward tangent sector for smooth manifolds with kinks. This is analogous to the inward tangent sector of Euclidean domains as discussed in Section 2.2. Our definition builds upon the classical characterization of the tangent space in terms of initial velocities of curves, which holds true for manifolds without/with boundary (see e.g. [21, p. 68–70]). For manifolds with boundary, one can identify inward tangent vectors at boundary points by specifying the domain of the curves we choose. This is how we came up with the following natural definition.